10. (8 marks) (using dataset: meap93, in R: data(meap93, package-wooldridge)) We want to explore the relationship betweenI need to know the code to do this question in R

thanks
Answer #1

i) the population model value for sci11, when totcomp =tc  is

\text{sci11}(tc)=\beta_0+\beta_1\log(\text{tc})+u

The value of sci11 when totcomp is increased by 10%, that is when totcomp=tc*1.1 is

= ß0 + β1 (log(tc) log( 1.1)) + u using the fact that log(ab) = log(a) o Bi log(tc)1 log(1.1) -scil1 (tcB1 x 0.095 log(b)

That means for a 10% increase in totcomp, the percentage point change in sci11 is

\begin{align*} \beta_1\times 0.095\approx \beta_1/10 \end{align*}

ii) The R code to estimate the model is below (all statements starting with # are comments and can be removed)

#install the package for the first time
install.packages('wooldridge')

#Load the data meap93
data(meap93,package='wooldridge')
#print some records
head(meap93)

#estimate the regression equation
fit<-lm(sci11~log(totcomp),data=meap93)
summary(fit)

# get this output

Call: lm (formula = sci II ~ log (totcomp), data = meap93) Residuals: 10 Median 3Q Max Min 42.991 -7.810 -0.2507.783 39.534 C

the estimated equation is

-69.771 + 11.286 log(totcomp) scil 1

R20.01858

Sample size n=408

iii) from part i) we can say that if the total compensation increases by 10%, then the estimated percentage point increase in scill is

R code

inc<-log(1.1)*fit$coef[2]
sprintf('The increase in scill is %.2f',inc)

##output is

> sprint f( The increase in scill is %.2f, inc) [11 The increase in scill is 1.08 13 I

Or using the approximation from part i) the increase is

11,286/10 = 1.13 percentage points

ans: If compensation increases by 10% the estimated percentage point increase in sci11 is 1.1

iv) Solving for totcomp using the fitted model we get

sci-69.771 11.286 log(totcomp) log(totcomp)- SCi11 +69.771 11.286 scil1+69.771 totcomp e 11.286

For the fitted values of sci11 to be greater than 100, we need the

-3, 411,343 totcomp 112863

That is the total compensation for the teachers to be more than $3,411,343, which is very unlikely for a school teacher in the US to aspire to. The maximum total compensation in the dataset is $63,518.

max (meap93stotcomp) 1 63518

Hence it is unlikely that in this dataset corresponding to the salary of school teachers, the fitted values of sci11 can be greater than 100.